OpenStudy (anonymous):

x^2-2kx+49 is the square of a binomial. What is the possible value of k?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

how do you solve that? Can you show your work pleazzz???

OpenStudy (anonymous):

(x+a)^2=x^2+2ax+a^2=x^2-2kx+49 a^2=49 2ax=-2kx a=+/- 7 a=-k k=-a k=+/- 7

OpenStudy (anonymous):

perfect square \[(a-b)^2=a^2-2ab+b^2\] here \[a=x\] \[b^2=49\] so \[b=7\]

OpenStudy (anonymous):

wait rsvitale, where do you get the a from?

OpenStudy (anonymous):

k could be -7 too, (x+7)^2=x^2+14x+49, which would mean k=-7

OpenStudy (anonymous):

a is just a variable for the second term in the binomial

OpenStudy (anonymous):

why is it (a-b)^2 im confused about the beginning of the problem

OpenStudy (anonymous):

thats the general form of a square of a binomial

OpenStudy (anonymous):

so we want to see what the general form looks like expanded and find what k can be

OpenStudy (anonymous):

\[(a+b)^2=(a+b)(a+b)=a^2+ab+ba+b^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

ok satelite, now explain from there. i get that step now so what do we do next... explain it like you just did.... sorry im not very good at math